OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. Q(3, 5), m = no slope

OpenStudy (anonymous):

@hartnn Can you help me?

hartnn (hartnn):

the equation of line through (x1,y1) and slope m is (y-y1)=m(x-x1)

OpenStudy (anonymous):

I've got y-5= ? (x-3) What do I do when there is no slope? Do I leave it blank?

hartnn (hartnn):

no slope means n= infinity, which means its a vertical line . vertical line have the form of x=c so here what is x value?

OpenStudy (anonymous):

3. So the slope would be the equivalent of 3 in this equation?

hartnn (hartnn):

no, no slope is infinite, x=3 is the equation of the line.

OpenStudy (anonymous):

Oh, I get it now! Thanks!

hartnn (hartnn):

welcome ^_^